OpenStudy (anonymous):

Please help! I only have 20 more minutes! Sergio has two similar cylindrical pots, Pot A and Pot B. The radius of Pot A is 20 inches, and the radius of Pot B is 4 inches. What is the ratio of the volume of Pot A to the volume of Pot B?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

For similar figures, ratio of volumes = (ratio of lengths)^3

ganeshie8 (ganeshie8):

so simply find the ratio of radii, then cube it

OpenStudy (anonymous):

Okay well for Pot A the formula would be: V= 3.14 x 20^2 x h. Right? Which then would give you: V = 3.14 x 400 x h?

ganeshie8 (ganeshie8):

we don't really need to all that

ganeshie8 (ganeshie8):

`The radius of Pot A is 20 inches, and the radius of Pot B is 4 inches. ` ratio of radii = 20/4 = 5 therefore, ratio of volumes = 5^3 = 125

OpenStudy (anonymous):

Was that my answer?

OpenStudy (anonymous):

I guess so.